please help me slove this question f(x)=x^2+7
i got this but i dont know what to do from here -7+ or - the square root of 85 over 2

Answer by Alan3354(69443)

f(x)=x^2+7
x^2 + 7 = 0
x^2 = -7
x = ± isqrt(7) (i = sqrt(-1) )